# Charge and spin current pumping by ultrafast demagnetization dynamics

### v1

@@ Abstract
Conclusions and Outlook.—By using time dependence [Fig. 1] of ultrafast demagnetization [Eq. (1)] from experiments [7] on fsLP-driven Ni layer—which is plugged into a two-terminal setup [Fig. 2] of standard theory [24, 25] of spin pumping to replace its slowly and harmonically precessing magnetization of fixed length driven by microwave absorption [35]—we directly connect these two apparently disparate phenomena. That is, time-dependent quantum transport theory that can handle [28, 29, 30, 31] arbitrary time-dependence of LMMs within the central FM region in Fig. 2 shows how demagnetization pumps, surprisingly, both spin and charge currents. The physical picture emerging is that of fsLP which drives electrons far from equilibrium to cause their currents oscillating at light frequency (as well as at its high harmonics [22, 2]), while the ensuing [6] demagnetization dynamics causes additional charge currents whose time derivative [Eq. (5)] generates spectral features of emitted EM radiation at THz frequencies [Fig. 4(c)]. We emphasize that slow recovery [Fig. 1] of magnetization on longer $\sim 1$ ps timescales does not generate any EM radiation with features in 0.1–30 THz range [Fig. 4(d)]. Our prediction of direct charge pumping by ultrafast demagnetization dynamics offers a unified explanation for experimentally observed THz radiation from both single FM layer [10] and FM/NM bilayers [12, 13, 14, 20]. It can be easily confirmed or falsified by observing THz radiation from, e.g., Pt/Ni/Pt trilayer whose intensity is comparable to the one from often employed [20] Ni/Pt bilayer. In contrast, standard phenomenological picture [12, 13, 14, 20] of spin-to-charge current conversion by ISHE in Pt would predict no EM radiation from Pt/Ni/Pt trilayer due to opposite direction of ISHE charge currents within two Pt layers. Even if much smaller THz radiation is found in Pt/Ni/Pt trilayer, our theory still provides microscopic explanation for the origin of spin current flowing from Ni to Pt layer, thereby replacing phenomenologically conjectured spin voltage [23, 14] as its driving mechanism.

### v2

@@ Abstract
Conclusions and Outlook.—By using time dependence [Fig. 1] of ultrafast demagnetization [Eq. (1)] from experiments [7] on fsLP-driven Ni layer—which is plugged into a two-terminal setup [Fig. 2] of standard theory [24, 25] of spin pumping to replace its slowly and harmonically precessing magnetization of fixed length driven by microwave absorption [35]—we directly connect these two apparently disparate phenomena. That is, time-dependent quantum transport theory that can handle [28, 29, 30, 31] arbitrary time-dependence of LMMs within the central FM region in Fig. 2 shows how demagnetization pumps, surprisingly, both spin and charge currents. The physical picture emerging is that of fsLP which drives electrons far from equilibrium to cause their current oscillating at light frequency (as well as at high harmonics of light frequency [22, 2]), while the ensuing [6] demagnetization dynamics pumps additional charge current whose time derivative [Eq. (5)] generates spectral features of emitted EM radiation at THz frequencies in Fig. 4(c) [additional Fig. S1 in the Supplemental Material  [^1] See Supplemental Material at https://wiki.physics.udel.edu/qttg/Publications, which includes Ref. [54], for additional Fig. S1 which plots orange curves from Figs. 4(a) and  4(c), obtained when only demagnetization dynamics drives electrons out of equilibrium, as well as their two counterpart where both fsLP and demagnetization dynamics or just fsLP are present as generators of electronic currents. demonstrates that presence of both fsLP and demagnetization dynamics as concurrent nonequilibirum drives does not change our conclusions from Fig. 4(c)]. We emphasize that slow recovery [Fig. 1] of magnetization on longer $\sim 1$ ps timescales does not generate any EM radiation with features in 0.1–30 THz range [Fig. 4(d)]. Our prediction of direct charge pumping by ultrafast demagnetization dynamics offers a unified explanation for experimentally observed THz radiation from both single FM layer [10] and FM/NM bilayers [12, 13, 14, 20]. It can be easily confirmed or falsified by observing THz radiation from, e.g., Pt/Ni/Pt trilayer whose intensity is comparable to the one from often employed [20] Ni/Pt bilayer. In contrast, standard phenomenological picture [12, 13, 14, 20] of spin-to-charge current conversion by ISHE in Pt would predict no EM radiation from Pt/Ni/Pt trilayer due to opposite direction of ISHE charge currents within two Pt layers. Even if much smaller THz radiation is found in Pt/Ni/Pt trilayer, our theory still provides microscopic explanation for the origin of spin current flowing from Ni to Pt layer, thereby replacing phenomenologically conjectured spin voltage [23, 14] as its driving mechanism.